OpenStudy (anonymous):

simplify: -24x^6y^7 over 18x^-3y^9

OpenStudy (turingtest):

use the following formula:\[\large\frac{x^a}{x^b}=x^{a-b}\]

OpenStudy (anonymous):

what is a and what is b?

OpenStudy (turingtest):

whatever the exponents happen to be

OpenStudy (turingtest):

examples\[\frac{x^3}{x^2}=x^{3-2}=x\]\[\frac{x^2}{x^5}=x^{2-5}=x^{-3}\]\[\frac{x^2}{x^{-3}}=x^{2-(-3)}=x^5\]